# Approximate Nearest Neighbour Search on Dynamic Datasets: An Investigation

### v2

@@ Abstract
Approximate k-Nearest Neighbour (ANN) methods are often used for mining information and aiding machine learning on large scale high-dimensional datasets. ANN methods typically differ in the index structure used for accelerating searches, resulting in various recall/runtime trade-off points. For applications with static datasets, runtime constraints and dataset properties can be used to empirically select an ANN method with suitable operating characteristics. However, for applications with dynamic datasets, which are subject to frequent online changes (like addition of new samples), there is currently no consensus as to which ANN methods are most suitable. Traditional evaluation approaches do not consider the computational costs of updating the index structure, as well as the frequency and size of index updates. To address this, we empirically evaluate 5 popular ANN methods on two main applications (online data collection and online feature learning) while taking into account these considerations. Two dynamic datasets are used, derived from the SIFT1M dataset with 1 million samples and the DEEP1B dataset with 1 billion samples. The results indicate that the often used k-d trees method is not suitable on dynamic datasets as it is slower than a straightforward baseline exhaustive search method. For online data collection, the Hierarchical Navigable Small World Graphs method achieves a consistent speedup over baseline across a wide range of recall rates. For online feature learning, the Scalable Nearest Neighbours method is faster than baseline for recall rates below 75%.
1 Introduction
@@
A notable limitation of traditional ANN evaluation approaches is the assumption that the index construction is a one-off offline process, as shown in Fig. 1. In turn this means that traditional approaches do not take into account the computational costs of recomputing/updating the index structure for dynamic datasets, and do not consider the frequency and size of index updates. Hence there is currently no consensus as to which ANN methods are most suitable for use on dynamic datasets.
To address this shortcoming, in this work we evaluate111Associated source code available at https://github.com/data61/DyANN the appropriateness of 5 popular ANN methods while explicitly taking into account the characteristics of search on dynamic datasets, which includes the costs of index updates as well as the rate and magnitude of the updates. The evaluations are performed on two dynamic datasets that are representative of broad application areas. The first dataset is modelled after real-time applications, where each subsequent query is appended as a new sample to the dataset. The second dataset represents online machine learning applications, where samples are converging across incremental updates.
We continue the paper as follows. Section 2 briefly overviews the main families of ANN search methods. Section 3 covers the characteristics of search on dynamic datasets. Section 4 empirically evaluates the covered ANN approaches on two dynamic datasets. Concluding remarks and possible avenues for further research are given in Section 5.
2 ANN Search Families
@@
The above ANN methods have parameters with recommended ranges that are typically obtained and tuned using static search problems [2, 17]. As the nature of search on dynamic datasets is notably different (see Section 3), such recommended parameter ranges may not align with the optimal range for dynamic search problems.
3 Characteristics of Search on Dynamic Datasets
We focus on two main application scenarios on dynamic datasets: (i) online data collection, and (ii) online feature learning. Progressively growing a dataset is an example of online data collection. Within many machine learning (ML) processes, the training of an embedding space [8, 10] is an example of online feature learning. Updating ML model parameters during the learning process will update the embedded representation of indexed samples. This update can affect ANN index structures and hence degrade the performance of subsequent searches.
@@ 3 Characteristics of Search on Dynamic Datasets
The batching of events and searches can also impact the trade-off between accuracy and runtime in a less direct way. For a fixed amount of compute in a given period of time, let us consider a fixed number of events and searches that will be processed. At each extreme of batching we can have events and searches alternating one after another, or we can have a large chunk of events clumped together followed by a large chunk of the searches. Examples are an autonomous vehicle capturing and querying for each location as it travels, and a closed loop learning pipeline completing a full forward pass to then feeding back neighbourhood information about an entire dataset [21].
@@
4 Experiments
We perform a comparative evaluation of the ANN methods overviewed in Section 2 on two dynamic datasets (described below) representing online data collection and online feature learning. All methods were limited to run on a single thread on AMD EPYC 7543 CPU, using up to 10 GB RAM. Restricting execution to a single thread aims to mitigate potential performance biases that can occur at particular parallelisation thresholds.
@@ 4 Experiments
4.1 Datasets
The Online Data Collection (ODC) dataset is modelled after real-time applications where each subsequent query is appended as a new sample to the dataset. We use various subsets of the SIFT1M dataset [1, 13] which contains 1 million samples with 128 dimensions. We use addition events (defined in Section 3) to grow each subset to double its initial size. Initial samples are expected to be representative of the total sample pool, and altering this may degrade the performance of some methods.
The Online Feature Learning (OFL) dataset represents a generalised but flexible model for machine learning applications. We apply update events (defined in Section 3) to model a dataset where samples are converging across incremental updates. Various subsets of the DEEP1B dataset [3] are used, which contains 1 billion samples with 96 dimensions.
4.2 Comparative Evaluation
@@
The results are shown in Figs. 2 and 3 for the ODC and OFL datasets, respectively, in terms of speedup over brute-force search across a range of accuracy points. Accuracy is measured in terms of average recall, where each search returns the top-50 nearest neighbours to a query and is scored against the ground truth top-50. All neighbours are weighted equally for scoring. This choice of neighbourhood size and weighting aims to approximate real-world application scenarios. In addition to search runtime, the computation time for each method also includes event processing time (defined in Section 3).
@@
4.3 Effect of Update Event Rates and Batch Sizes
Update events can be quite disruptive in that they can affect all samples in an index. To gauge the importance of update event rates as well as batch sizes (defined in Section 3) we perform experiments where are these aspects are varied. For this task we use the HNSWG method on the OFL dataset, as the results from Section 4.2 show that the method is capable of operating over a wide range of recall points. We evaluate two scenarios: (a) increasing update event rate with fixed search rate, and (b) increasing update event batch size with fixed search batch size.
@@ 5 Concluding Remarks
Future avenues of research include studying the effects of pruning a dataset (eg. removal of old or low saliency samples), which is an important consideration for bounding memory use. Furthermore, it would be useful to evaluate the effects of temporal dependence (mutual information) between consecutive samples due to temporal consistency in applications such as robotic localisation [7]. Exploiting mutual information between consecutive search batches may provide large performance gains, due to the possibility of basing the starting point of each search on the search results obtained for preceding samples.

### v4

@@ Abstract
Approximate k-Nearest Neighbour (ANN) methods are often used for mining information and aiding machine learning on large scale high-dimensional datasets. ANN methods typically differ in the index structure used for accelerating searches, resulting in various recall/runtime trade-off points. For applications with static datasets, runtime constraints and dataset properties can be used to empirically select an ANN method with suitable operating characteristics. However, for applications with dynamic datasets, which are subject to frequent online changes (like addition of new samples), there is currently no consensus as to which ANN methods are most suitable. Traditional evaluation approaches do not consider the computational costs of updating the index structure, as well as the rate and size of index updates. To address this, we empirically evaluate 5 popular ANN methods on two main applications (online data collection and online feature learning) while taking into account these considerations. Two dynamic datasets are used, derived from the SIFT1M dataset with 1 million samples and the DEEP1B dataset with 1 billion samples. The results indicate that the often used k-d trees method is not suitable on dynamic datasets as it is slower than a straightforward baseline exhaustive search method. For online data collection, the Hierarchical Navigable Small World Graphs method achieves a consistent speedup over baseline across a wide range of recall rates. For online feature learning, the Scalable Nearest Neighbours method is faster than baseline for recall rates below 75%.
I Introduction
@@
A notable limitation of traditional ANN evaluation approaches is the assumption that the index construction is a one-off offline process, as shown in Fig. 1. In turn this means that traditional approaches do not take into account the computational costs of recomputing/updating the index structure for dynamic datasets, and do not consider the rate and size of index updates. Hence there is currently no consensus as to which ANN methods are most suitable for use on dynamic datasets.
To address this shortcoming, in this work we evaluate111 Associated source code is available at https://github.com/data61/DyANN. The code is extensible with new ANN methods and new categories of dynamic search problems. the appropriateness of 5 popular ANN methods while explicitly taking into account the characteristics of search on dynamic datasets, which includes the costs of index updates as well as the rate and magnitude of the updates. The evaluations are performed on two dynamic datasets that are representative of broad application areas. The first dataset is modelled after real-time applications, where each subsequent query is appended as a new sample to the dataset. The second dataset represents online machine learning applications, where samples are converging across incremental updates.
We continue the paper as follows. Section II briefly overviews the main families of ANN search methods. Section III covers the characteristics of search on dynamic datasets. Section IV empirically evaluates the covered ANN approaches on two dynamic datasets. Concluding remarks and possible avenues for further research are given in Section V.
II ANN Search Families
@@
The above ANN methods have parameters with recommended ranges that are typically obtained and tuned using static search problems [2, 17]. As the nature of search on dynamic datasets is notably different (see Section III), such recommended parameter ranges may not align with the optimal range for dynamic search problems.
III Characteristics of Search on Dynamic Datasets
We focus on two main application scenarios on dynamic datasets: (i) online data collection, and (ii) online feature learning. In online data collection, a dataset is progressively growing via addition of new samples. Often the new samples are compared against existing samples in order to achieve detection of anomalies or estimation of saliency. Within many machine learning (ML) processes, the training of an embedding space [8, 10] is an example of online feature learning. Updating ML model parameters during the learning process will update the embedded representation of indexed samples. This update can affect ANN index structures and hence degrade the performance of subsequent searches.
@@ III Characteristics of Search on Dynamic Datasets
The batching of events and searches can also impact the trade-off between accuracy and runtime in a less direct way. For a fixed amount of compute in a given period of time, let us consider a fixed number of events and searches that will be processed. At each extreme of batching we can have events and searches alternating one after another (eg. an autonomous vehicle capturing and querying for each location as it travels), or we can have a large chunk of events clumped together followed by a large chunk of the searches (eg. a closed loop learning pipeline completing a full forward pass to then feeding back neighbourhood information about an entire dataset [22]).
@@
IV Experiments
We perform a comparative evaluation of the ANN methods overviewed in Section II on two dynamic datasets (described below) representing online data collection and online feature learning. All methods were limited to run on a single thread on AMD EPYC 7543 CPU, using up to 10 GB RAM. Restricting execution to a single thread aims to mitigate potential performance biases that can occur at particular parallelisation thresholds.
@@ IV Experiments
IV-A Datasets
The Online Data Collection (ODC) dataset is modelled after real-time applications where each subsequent query is appended as a new sample to the dataset. We use various subsets of the SIFT1M dataset [1, 13] which contains 1 million samples with 128 dimensions. We use addition events (defined in Section III) to grow each subset to double its initial size. Initial samples are expected to be representative of the total sample pool, and altering this may degrade the performance of some methods.
The Online Feature Learning (OFL) dataset represents a generalised but flexible model for machine learning applications. We apply update events (defined in Section III) to model a dataset where samples are converging across incremental updates. Various subsets of the DEEP1B dataset [3] are used, which contains 1 billion samples with 96 dimensions.
IV-B Comparative Evaluation
@@
The results are shown in Figs. 2 and 3 for the ODC and OFL datasets, respectively, in terms of speedup over brute-force search across a range of accuracy points. Accuracy is measured in terms of average recall, where each search returns the top-50 nearest neighbours to a query and is scored against the ground truth top-50. All neighbours are weighted equally for scoring. This choice of neighbourhood size and weighting aims to approximate real-world application scenarios. In addition to search runtime, the computation time for each method also includes event processing time (defined in Section III).
@@
IV-C Effect of Update Event Rates and Batch Sizes
Update events can be quite disruptive in that they can affect all samples in an index. To gauge the importance of update event rates as well as batch sizes (defined in Section III) we perform experiments where are these aspects are varied. For this task we use the HNSWG method on the OFL dataset, as the results from Section IV-B show that the method is capable of operating over a wide range of recall points. We evaluate two scenarios: (a) increasing update event rate with fixed search rate, and (b) increasing update event batch size with fixed search batch size.
@@ V Concluding Remarks
Future avenues of research include studying the effects of pruning a dataset (eg. removal of old or low saliency samples), which is an important consideration for bounding memory use. Furthermore, it would be useful to evaluate the effects of temporal dependence (mutual information) between consecutive samples due to temporal consistency in applications such as robotic localisation [7]. Exploiting mutual information between consecutive search batches may provide large performance gains, due to the possibility of basing the starting point of each search on the search results obtained for preceding samples. Lastly, it may be beneficial to explore the trade-offs between accuracy and explainability of each ANN method, in order to take into account aspects of responsible AI [20].